Two identical lamps, each of resistance 8Ω are connected in parallel. They are connected to a 12V battery.
Use the equation I=v/r to calculate the current through one lamp
1
Expert's answer
2018-10-31T13:56:09-0400

According to the law ohm, the current in the circuit is

I=U/R

I=12/8=1.5A